# Scaling Laws for Black-box Adversarial Attacks

### v1

@@ Abstract
A longstanding problem of deep learning models is their vulnerability to adversarial examples, which are often generated by applying imperceptible perturbations to natural examples. Adversarial examples exhibit cross-model transferability, enabling to attack black-box models with limited information about their architectures and parameters. Model ensembling is an effective strategy to improve the transferability by attacking multiple surrogate models simultaneously. However, as prior studies usually adopt few models in the ensemble, there remains an open question of whether scaling the number of models can further improve black-box attacks. Inspired by the findings in large foundation models, we investigate the scaling laws of black-box adversarial attacks in this work. By analyzing the relationship between the number of surrogate models and transferability of adversarial examples, we conclude with clear scaling laws, emphasizing the potential of using more surrogate models to enhance adversarial transferability. Extensive experiments verify the claims on standard image classifiers, multimodal large language models, and even proprietary models like GPT-4o, demonstrating consistent scaling effects and impressive attack success rates with more surrogate models. Further studies by visualization indicate that scaled attacks bring better interpretability in semantics, indicating that the common features of models are captured.
@@ 1 Introduction
Over the past decade, deep learning has achieved remarkable advancements across various computer vision tasks, leading to its widespread application in practical contexts [52]. However, deep neural networks (DNNs) remain susceptible to adversarial examples [33, 76] – perturbations that are subtle and nearly imperceptible but can significantly mislead the model, resulting in incorrect predictions. Adversarial attack has garnered a considerable interest due to its implications for understanding DNN functionality [23], assessing model robustness [100], and informing the development of more secure algorithms [63].
@@
Figure 1: Key insights of our scaling laws for the transfer-based black-box attack regarding the number of surrogate models. By scaling the model ensembling, we observe that the attack success rate (ASR) and the discriminative loss on the target model follow the scaling laws, and these two metrics have a clear symmetry between each other. As the cardinality of models scales, the semantic interpretability of the adversarial perturbations keeps increasing by visualization. Details of our analyses are shown in Sec. 4.
@@
Adversarial attacks are typically categorized into white-box and black-box attacks. Given the greater risks with commercial services, e.g., GPT-4 [66], black-box attack methodologies [54, 24] have become a growing focus. One prominent approach within black-box attacks is the transfer-based attack, which approximately optimizes the adversarial examples against the black-box model by minimizing the discriminative loss on a selection of predefined surrogate models. Researchers have devised advanced optimization algorithms [63, 85, 54] and data augmentation techniques [61, 25] in the context of transfer attacks. Model ensembling [58, 24] further facilitates these methods by leveraging multiple surrogate models simultaneously to improve the robustness of adversarial examples. By averaging the gradients or predictions from different models, ensemble attack methods seek to increase the transferability across various target models [24, 13].
Despite the progress in ensemble attack algorithms, the number of surrogate models used is often limited, leaving the concept of model ensemble scaling under-explored. In contrast to prior works, we hereby investigate the scaling laws of ensemble attacks. Notably, the concept of scaling laws has led to substantial advancements in pretrained models, exemplified by the development of the Generative Pretrained Transformer (GPT) [46]. Traditional scaling laws aim for better generalization to unseen scenarios by training on massive data, while ensemble attacks perform optimization on multiple models for better transferability. These two claims are indeed equivalent as prior works [24, 13] have pointed out that the transferability of adversarial examples generated from surrogate models parallels the generalization observed in neural network training. This equivalence prompts an intriguing question: Can we achieve significant improvements in transfer-based attack by scaling the number of surrogate models, akin to the scaling of data and model parameters in GPT?
In this study, we establish the scaling laws of black-box ensemble attacks regarding the number of surrogate models. We first perform a theoretical analysis on the convergence of ensemble attacks, which motivates our further empirical studies. By conducting extensive experiments of targeted attacks, we systematically explore the scaling of model ensembling, uncovering clear trends that inform our understanding of scaling behavior in ensemble attacks. Besides near 100% attack success rates against black-box classifiers, we also frequently mislead the predictions of modern Vision Language Models (VLMs) to predefined objects, solely with a set of pretrained classification models.
Beyond these studies on open-source models, we leverage our insights to conduct scaled attacks on proprietary models like GPT-4 [66]. We extend the attack methods to a group of CLIP vision encoders to craft targeted adversarial examples, which achieve a success rate of 90% against GPT-4o and GPT-4-Turbo, and over 60% against other commercial MLLMs. This highlights the vulnerabilities of these models to perturbations and the significance of the scaling laws for ensemble attack. We also explain the existence of the scaling laws in ensemble attacks from two aspects via further studies. The increasing semantic interpretability of the visualized perturbations suggests that the adversarial examples capture the common features of models on the natural image manifold, while the loss of effectiveness on adversarially trained models indicates that the scaling comes from the shareable feature space across diverse models [90].
2 Related Work
2.1 Black-Box Attacks
2.2 Ensemble Attacks
Prior studies [39] indicate that increasing the number of classifiers in adversarial attacks can effectively reduce the upper bound of generalization error in Empirical Risk Minimization (ERM). Recent research [97] has theoretically defined transferability error and sought to minimize it through model ensemble. Nonetheless, while an upper bound is given, they do not really scale up in practice and have to train neural networks from scratch to meet the assumption of sufficient diversity among ensemble models [57].
2.3 Scaling Laws
In the domain of adversarial attack research, recent work has attempted to derive scaling laws for adversarially trained defense models on CIFAR-10 [48, 8]. However, the broader concept of scaling remains underexplored. While some preliminary observations on the effect of enlarging model ensemble can be found in prior studies [97, 3], these effects have not been systematically analyzed. In this paper, however, we focus on the empirical scaling laws within model ensembles, and verify it through extensive experiments on open-source pretrained models. Motivated by this perspective, we unveil vulnerabilities within commercial LLMs by targeting their vision encoders.
@@ 3 Backgrounds
In this section, we propose necessary backgrounds including problem formulation and algorithms.
@@ 3.2 Prerequisite Algorithms
Researchers have been developing sophisticated algorithms for ensemble attacks, and we select SOTA algorithms from them to conduct our experiments.
Momentum Iterative Fast Gradient Sign Method. MI-FGSM [24] stabilizes the gradient updating directions by integrating a momentum term into the iterative process, which generates the momentum by
@@
and updates $\bm{x}$ by
$\bm{x}_{t+1}=\bm{x}_{t}+\alpha\cdot\operatorname{sign}(\bm{g}_{t+1}).$ (4)
Common Weakness Method. The common weakness attack (CWA) [13] seeks to identify shared vulnerabilities across surrogate models by encouraging flatter loss landscapes and closer proximity between local optima of these models, which is minimizing the second-order Taylor series
@@
Figure 2: Comparisons between CWA and MI-FGSM methods. We measure the averaged gradient magnitudes on $\bm{x}$ after 10 steps of iteration, and record the results across $T$ from $1$ to $30$. On the right side there are the visualizations of gradients of $T=30$.
CWA enables the scaling laws. Since the gradients of normally trained models commonly lack semantics and have random directions, naive ensemble methods that directly average these gradients lead to strong gradient conflicts, significantly impacting optimization. As shown in Fig. 2, as the number of models increases, gradient conflicts among other optimizers (MI-FGSM [24] here) become more severe, eventually reducing the ensemble gradient to zero and resulting in complete optimization stagnation. CWA not only fully avoids this issue but also promotes gradient alignment, causing the gradient magnitude to increase as the number of models grows. This demonstrates that CWA achieves a significantly higher information utilization rate than other optimizers, avoiding gradient conflict, and thus enabling the scaling laws.
Spectrum Simulation Method. The spectrum simulation attack (SSA) [61] enhances adversarial transferability by applying a spectrum transformation $\mathcal{T}$ to the input data and calculate gradients with respective to $\mathcal{T}(\bm{x})$ instead of $\bm{x}$. As the variable of differentiation changes, SSA cannot combine with MI, but both of them can combine with CWA by changing their optimization objectives. As shown in Table 1, SSA achieves the success rate of 10% higher compared with MI algorithm, demonstrating its effectiveness of generating diverse spectrum saliency maps [61]. This encourages us to select it as key algorithm in our experiments.
ASR(%) Target: Car Target: Ship
@@
ResNet50 0.69 0.79 0.66 0.77
Table 1: Comparison of ASR between SSA and MI algorithms. We follow the experiment settings in Sec. 4.2, and measure the ASR on 2 target classes and 3 black-box models.
@@ 4 Scaling Laws for Adversarial Attacks
In this section, we bring our key findings on scaling laws for black-box adversarial attacks. Sec. 4.1 introduces fundamental concepts and theorems that give us motivations for the study of scaling laws. Sec. 4.2 contains our experiment settings and derivations of our empirical scaling laws. Sec. 4.3 demonstrates the results when we take vision language models as black-box models.
@@ 4.1 Motivations
Firstly, we observe that in most practical scenarios, $\bm{x}_{\text{nat}}$ and $\bm{y}$ are predefined and fixed. Consequently, Eq. (1) and Eq. (2) can be considered as functions of variable $\bm{x}$ only.
@@
Define $\bm{x}^{*}$ and $\hat{\bm{x}}$ to be the minimizer of Eq. (1) and Eq. (2) respectively, which can be written as
@@
These definitions are used to simplify our notations and better demonstrate our proofs. Then we bring our key observations that are shown in the following theorem.
@@
When we replace $\mathcal{L}$ with cross-entropy loss in Theorem 1, we get Theorem 2 by some simple derivations.
@@
Following the notation and assumptions in Theorem 1, when using negative likelihood loss (cross entropy), we have:
@@
Theorem 2 gives us motivations on further exploring the affects of ensemble cardinality $T$ as empirical loss converges with a factor of $T$. This leads us to recovering the scaling laws on black-box adversarial attacks.
@@ 4.2 Scaling Laws on Image Classifiers
Surrogate Models. To ensure the transferability and diversity of models, we select 64 models, with 62 pretrained models from Torchvision [64] and Timm [87] with different hyperparameters, architectures, and training datasets – Alexnet [49], Densenet [38], Efficientnet [79], Googlenet [77], Inception-V3 [78], Maxvit [84], Mnasnet [80], Mobilenet [37], Regnet [96], Resnet [35], Shufflenet [99], Squeezenet [41], Vgg [73], ViT [28], WideResnet [98], Beit [7], Ghostnet [34], Deit [81], LCnet [17, 88], Repvgg [22], Dpn [14], Mit [93], Bit [47], CvT [91], and 2 adversarially trained models available on Robustbench - FGSMAT [50] with Inception-V3, Ensemble AT (EnsAT) [83] with Inception-ResNet-V2. The full list of our surrogate models can be found in Appendix B.
Figure 3: Scaling Laws over model ensemble in a black-box setting. We plot the x-axis on a base-2 logarithmic scale, and we measure both the attack success rate (ASR) and averaged cross entropy loss of test models. The cardinality of the model ensembles is varied from $2^{0}$ to $2^{6}$, with models randomly selected for each ensemble. We fit the results of 4 test models by lines for better demonstration.
Black-box Models. We carefully choose models that are both widely known and are different from training ensembles in architecture and hyperparameters. We select 7 models as our target models, including Swin-Transformers [59], ConvNeXt [60], and ResNeXt [94].
Dataset. Following previous works [45], we randomly take 100 images from the NIPS17 dataset ***https://www.kaggle.com/competitions/nips-2017-non-targeted-adversarial-attack. All images are resized into 224 $\times$ 224. We select 5 categories as target labels, including cat, car, ship, deer and bird.
Hyperparameters. We set the perturbation budget $\epsilon=8/255$ under $\ell_{\infty}$ norm. We take 40 iteration steps of SSA-CWA, while other hyper-parameter settings are the same as in [13]. We set the loss function $\mathcal{L}$ to be cross entropy loss between the output logit $f(\bm{x})$ and the label $\bm{y}$.
@@
  • For a fixed adversarial target, $\alpha$ depends solely on the model architecture. Models with the same architecture share the same $\alpha$, regardless of hyperparameter settings.
@@
  • ASR acts as a good representation of the empirical loss $\mathbb{L}$, with low evaluation loss resulting in high ASR generally. In fact it can be observed from Fig. 3 that there exists strong symmetry between $\mathbb{L}$ and ASR.
@@
Figure 4: Scaling Laws on VLMs. We measure the ASR of test VLMs and fit the results with lines. Note that scaling laws still exists except for the fact that results vary largely between models and targets.
Our experiment results clearly show that transferability of adversarial images exists, independent of specific test models or adversarial targets. It is important to note that both our training and evaluation models are open-sourced and pretrained on ImageNet [20], which means they do not strictly satisfy the i.i.d. assumption discussed in Sec. 4.1. Therefore, the theorems derived in Sec. 4.1 should be viewed as a sign of convergence and are not directly related with the empirical result measured by our experiments. Another key improvement observed from our scaling laws is that the evaluation loss $\mathbb{L}$ decreases exponentially with the ensemble cardinality $T$, which is much tighter compared to bounds derived in previous works [39, 97].
4.3 Scaling Laws on Vision Language Models
Vision language models (VLMs) have demonstrated impressive abilities in image understanding, describing and reasoning [9, 70, 82]. In this part we show that our scaling laws persist in these VLMs.
Black-box Models. We still leverage the adversarial examples generated against traditional image classifiers and select 3 popular VLMs – LLaVA-v1.5-7b [56], DeepSeek-VL-7b [19], and Phi-3-vision [1] as test models.
Evaluation Metrics. As VLMs output prompts rather than logits, we adopt the definition of ASR from MultiTrust [100], which leverages GPT4 for determining whether the targeted object is stated in the description. In addition to assessing the correctness of the image classification, it introduces prompts to evaluate the degree of distortion in the crafted image. If GPT determines that the image is highly corrupted, the attack is considered unsuccessful, even if the test model misclassifies the image. This metric ensures our judgments of successful attacks are of high quality.

### v2

@@ Abstract
Adversarial examples usually exhibit good cross-model transferability, enabling attacks on black-box models with limited information about their architectures and parameters, which are highly threatening in commercial black-box scenarios. Model ensembling is an effective strategy to improve the transferability of adversarial examples by attacking multiple surrogate models. However, since prior studies usually adopt few models in the ensemble, there remains an open question of whether scaling the number of models can further improve black-box attacks. Inspired by the scaling law of large foundation models, we investigate the scaling laws of black-box adversarial attacks in this work. Through theoretical analysis and empirical evaluations, we conclude with clear scaling laws that using more surrogate models enhances adversarial transferability. Comprehensive experiments verify the claims on standard image classifiers, diverse defended models and multimodal large language models using various adversarial attack methods. Specifically, by scaling laws, we achieve 90%+ transfer attack success rate on even proprietary models like GPT-4o. Further visualization indicates that scaling laws exists on the interpretability and semantics of adversarial perturbations.
@@ 1 Introduction
Over the past decade, deep learning has achieved remarkable advancements across various computer vision tasks, leading to its widespread applications [54]. However, deep neural networks (DNNs) remain susceptible to adversarial examples [35, 73] – perturbations that are subtle and nearly imperceptible but can significantly mislead the model, resulting in incorrect predictions. Adversarial attacks have garnered considerable interest due to their implications for understanding DNN functionality [21], assessing model robustness [98], and informing the development of more secure algorithms [65].
@@
Figure 1: Key observation of scaling laws for the transfer-based black-box attack. By scaling the number of surrogate models in the ensemble, we observe that the attack success rate and the discriminative loss on the target model follow a clear scaling law.
@@
Adversarial attacks are typically categorized into white-box and black-box attacks based on whether the attacker has knowledge of the architecture and parameters of the target model. Given the greater risks with commercial services, e.g., GPT-4 [68], black-box attack methodologies [60, 22] have attracted growing attention. One prominent approach within black-box attacks is the transfer-based attack, which approximately optimizes the adversarial examples against the black-box model by minimizing the discriminative loss on a selection of predefined surrogate models. Researchers have devised advanced optimization algorithms [65, 81, 57] and data augmentation techniques [63, 23] in transfer-based attacks. Model ensembling [60, 22] further facilitates these methods by leveraging multiple surrogate models simultaneously to improve the transferability of adversarial examples. By averaging the gradients or predictions from different models, ensemble attack methods seek to increase the transferability across various target models [22, 9].
Despite the progress in ensemble attack algorithms, the number of surrogate models used in previous research is often limited [9, 97, 94]. Recent advancement in large language models shows that scaling training data can significantly improve the generalization ability [48]. Since the transfer-based attack is also a generalization problem that expects adversarial examples from surrogate models to generalize to target models, while the training data are surrogate models rather than text or images [22, 9], this equivalence prompts an intriguing question: Can we achieve significant improvements in transfer-based attacks by scaling the number of surrogate models, akin to the scaling of data and model parameters in large language models like GPT?
In this work, we investigate the scaling laws for black-box adversarial attacks. Theoretically, we provide an asymptotic analysis on the convergence rate of generation error and adversarial examples in Sec. 3.1. Empirically, we show that the attack success rate and the loss are converging exponentially. We verify this by comprehensive experiments over traditional image classifiers (Sec. 3.2), defended models including adversarial training [65, 86] and DiffPure [67] (Sec. 3.3), and modern multimodal large language models (Sec. 3.4 and Sec. 4) using various attack methods (Sec. 5). These experimental results demonstrate the universality of scaling laws in transfer-based attacks.
More concretely, through scaling laws, we achieve over 30% improvement against vision classifiers and over 10% improvement against defended models. By extending the surrogate models to a group of CLIP [70] vision encoders, we achieve over 90% attack success rate against GPT-4o and GPT-4-Turbo, and over 80% against other commercial MLLMs. This highlights the vulnerabilities of these models to perturbations and the significance of scaling laws for ensemble attacks. We also demonstrate clear scaling laws on the interpretability and semantics of adversarial perturbations. Leveraging our insights into scaling laws, we emphasize further research on evaluating black-box weaknesses of MLLMs by focusing on large-scale experiments, rather than designing intricate algorithms for a small set of surrogates, which may deepen the understanding of the weaknesses of large models, ultimately contributing to the development of more robust and trustworthy MLLMs.
@@ 2 Backgrounds
In this section, we provide necessary backgrounds, including problem formulation and algorithms.
@@ 2.2 Prerequisite Algorithms
Researchers have been developing sophisticated algorithms for ensemble attacks, and we select the state-of-the-art algorithms from them to conduct our experiments.
MI-FGSM [22] is a pioneering method of transfer-based attacks. It stabilizes the gradient update directions by integrating a momentum term into the iteration:
@@
and updates $\bm{x}$ by $\bm{x}_{t+1}=\bm{x}_{t}+\alpha\cdot\operatorname{sign}(\bm{g}_{t+1})$.
Common Weakness Method (CWA) [9] seeks to identify common vulnerabilities across surrogate models by encouraging flatter loss landscapes and closer proximity between the local optima of these models. It minimizes the second-order Taylor series:
@@
Spectrum Simulation Attack (SSA) [63] applies a spectrum transformation $\mathcal{T}$ to the $\bm{x}$ and calculate gradients with respective to $\mathcal{T}(\bm{x})$ instead of $\bm{x}$, and can be combined with CWA to form SSA-CWA algorithm (pseudocode is provided in Appendix C). As shown in Table 1, SSA-CWA gives the best attack success rate. Therefore, in this paper, we use SSA-CWA as the core algorithm. We also explore the scaling laws of other algorithms in Sec. 5.1.
ASR Target: Car Target: Ship
@@
ResNeXt50 0.69 0.79 0.66 0.77
Table 1: Comparison of attack success rate (ASR) between SSA and MI algorithms. We follow the experiment settings in Sec. 3.2, and measure the ASR on 2 target classes and 3 black-box models.
Figure 2: Comparisons between CWA and MI-FGSM methods. We draw inputs $\bm{x}$ randomly from the dataset in Sec. 3.2 and measure the averaged gradient magnitudes on $\bm{x}$ after 10 steps of iterations. The results are recorded across $T$ from $1$ to $30$. On the right side, there are the visualizations of gradients of $T=30$.
Mitigating Gradient Conflict. We discover a new phenomenon of gradient conflict when ensembling a large number of surrogates. As shown in Fig. 2, as the number of models increases, the conflict between gradients of different models using common optimizers (e.g., MI-FGSM) becomes more severe, eventually reducing the ensemble gradient to zero and resulting in complete optimization stagnation. We conjecture this is because the gradients of normally trained models commonly lack semantics and have random directions; thus, naïve ensemble methods that directly average these gradients lead to strong gradient conflicts, significantly impacting optimization. To address this problem, we suggest to use optimizers that reduce the variance of gradients (e.g., SVRE [92]) or encourage the cosine similarity of gradients (e.g., CWA [9]). As demonstrated in Fig. 2, these methods fully avoid this issue. We analyze the gradient conflict on different attackers in more detail in Sec. 5.1.
@@ 3 Scaling Laws for Adversarial Attacks
In this section, we present our key findings on scaling laws for black-box adversarial attacks. Sec. 3.1 introduces fundamental concepts and theorems that give us motivations for the study of scaling laws. Sec. 3.2 contains our experiment settings and derivations of empirical scaling laws. Sec. 3.3 demonstrates clear scaling laws when attacking common defended models. Sec. 3.4 shows the results over multimodal large language models.
@@ 3.1 Motivations
First, we observe that in most practical scenarios, $\bm{x}_{\text{nat}}$ and $\bm{y}$ are predefined and fixed. Consequently, Eq. (1) and Eq. (2) can be considered as functions of variable $\bm{x}$ only.
@@
Define $\bm{x}^{*}$ and $\hat{\bm{x}}$ to be the minimizers of Eq. (1) and Eq. (2), respectively, which can be written as
@@
These definitions are used to simplify our notations and better demonstrate our proofs. Then, we bring our key observations that are shown in the following theorem.
@@
When we replace $\mathcal{L}$ with the cross-entropy loss in Theorem 1, we obtain Theorem 2 by some simple derivations.
@@
Following the notation and assumptions in Theorem 1, when using the cross-entropy loss, we have:
@@
Theorem 2 gives us motivations for further exploring the effects of ensemble cardinality $T$ as empirical loss converges with a factor of $T$. This leads us to recovering the scaling laws on black-box adversarial attacks.
@@ 3.2 Scaling Laws on Image Classifiers
Figure 3: Scaling Laws over model ensemble in a black-box setting. We plot the x-axis on a base-2 logarithmic scale, and we measure both the attack success rate (ASR) and average cross entropy-loss of target models. The cardinality of the model ensembles is varied from $2^{0}$ to $2^{6}$, with models randomly selected for each ensemble. We fit the results of 4 target models by lines for better demonstration.
@@
Surrogate Models. To ensure the transferability and diversity of models, we select 64 models from Torchvision [66] and Timm [84] with different hyperparameters, architectures, and training datasets – Alexnet [51], Densenet [40], Efficientnet [76], Googlenet [74], Inception-V3 [75], Maxvit [80], Mnasnet [77], Mobilenet [39], Regnet [93], Resnet [37], Shufflenet [96], Squeezenet [43], Vgg [72], ViT [27], WideResnet [95], Beit [4], Ghostnet [36], Deit [78], LCnet [14, 85], Repvgg [20], Dpn [11], Mit [90], Bit [49], CvT [88], Twins-SVT [13], and NFNet [6]. The full list of our surrogate models can be found in Appendix B.
Target Models. We carefully choose models that are both widely known and are different from training ensembles in architecture and hyperparameters. We select 7 models as our target models, including Swin-Transformers [61], ConvNeXt [62], and ResNeXt [91].
Dataset. Following previous work [47], we randomly take 100 images from the NIPS17 dataset [^1] https://www.kaggle.com/competitions/nips-2017-non-targeted-adversarial-attack. All images are resized into 224 $\times$ 224. We select 5 categories as target labels, including cat, car, ship, deer and bird.
Hyperparameters. We set the perturbation budget $\epsilon=8/255$ under $\ell_{\infty}$ norm. We take 40 iteration steps of SSA-CWA, while other hyperparameter settings are the same as in [9]. We set the loss function $\mathcal{L}$ to be the cross-entropy loss between the output logit $f(\bm{x})$ and the label $\bm{y}$.
Metrics. We leverage attack success rate (ASR) as our metric, which is (adversarial examples that induce successful attacks /// all adversarial examples).
@@
  • For a fixed adversarial target, $\alpha$ depends solely on the model itself. Models with the same architecture share the same $\alpha$, regardless of hyperparameter settings.
@@
  • ASR acts as a good representation of the empirical loss $\mathbb{L}$ defined in Eq. (2). It can be observed from Fig. 3 that there exists strong symmetry between $\mathbb{L}$ and ASR.
@@
Our experiment results clearly demonstrate the universality of scaling laws, independent of specific test models or adversarial targets. It is important to note that both our training and evaluation models are open-sourced and pretrained on ImageNet [18], which means they do not strictly satisfy the i.i.d. assumption discussed in Sec. 3.1. Therefore, the theorems derived in Sec. 3.1 should be viewed as a sign of convergence and are not directly related with the empirical result measured by our experiments. Another key improvement observed from our scaling laws is that the evaluation loss $\mathbb{L}$ decreases exponentially with the ensemble cardinality $T$, which is much tighter compared to bounds derived in previous works [41, 94].
3.3 Scaling Laws on Defended Models
Figure 4: Scaling Laws for defended models. While our adversarial examples exhibit high attack success rates across defenses, clear linear scaling laws can be observed from the fitted lines.
Figure 5: Scaling Laws for MLLMs. We measure the ASR of target MLLMs and fit the results with lines. Note that scaling laws still exist, except for the fact that results vary largely between models and targets.
To protect neural networks from attacks, various defense mechanisms have been proposed. Our scaling laws work well with these defense mechanisms, showing that these laws tend to be a general result across defenses.
Target Models. We select six defended models as target models, including 2 adversarially trained models – FGSMAT [52] and Ensemble AT (EnsAT) [79], and 4 input purification defenses – high-level representation guided denoiser (HGD) [56], bit depth reduction (BDR) [24], JPEG compression [16], and DiffPure [67]. We use the same surrogate models and algorithms as Sec. 3.2.
Experiment Results. As demonstrated in Fig. 4, we observe a clear scaling law in all these defended models, further demonstrating its universality. By ensembling a large number of normally trained surrogate models in Sec. 3.2, we improve the attack success rate of DiffPure [67] by more than 30%, adversarially trained models by more than 20%, and high-level guided denoiser [56] by more than 70%. This demonstrates that simply increasing the number of surrogate models can break state-of-the-art defenses like DiffPure [67] and adversarial training [65].
3.4 Scaling Laws on MLLMs
Multimodal Large Language Models (MLLMs) [58, 68, 29] have demonstrated impressive abilities in vision understanding, captioning, and reasoning. In this section, we show that our scaling laws persist in these MLLMs.
Target Models. We still leverage the adversarial examples generated against traditional image classifiers and select 3 popular VLMs – LLaVA-v1.5-7b [58], DeepSeek-VL-7b [17], and Phi-3-vision [29] as target models.
Evaluation Metrics. As MLLMs output text rather than logits, we adopt the definition of ASR from MultiTrust [98], which leverages GPT-4 for determining whether the targeted object is stated in the description. In addition to assessing the correctness of the image classification, it introduces prompts to evaluate the degree of distortion in the crafted image. If GPT determines that the image is highly corrupted, the attack is considered unsuccessful, even if the target model misclassifies the image. This metric ensures that our judgments of successful attacks are of high quality.